OpenStudy (anonymous):

-3x^6y^4/x^-3y^-2 explain

OpenStudy (anonymous):

@Loser66

OpenStudy (loser66):

Again? hahaha.. just do the same. I'll check

OpenStudy (anonymous):

put them over each other as a fraction so i can look at it right lol

OpenStudy (loser66):

why?? you just write it on paper and see it

OpenStudy (anonymous):

well i dont have paper with me right now

OpenStudy (loser66):

\[\dfrac{-3x^6y^4}{x^{-3}y^{-2}}\]

OpenStudy (anonymous):

so -1x^2y^-2????

OpenStudy (loser66):

nope

OpenStudy (anonymous):

oh no its -3x^2y^2

OpenStudy (loser66):

nope again -3 is correct, but the exponents are wrong.

OpenStudy (anonymous):

-2 , -2???

OpenStudy (loser66):

6+3 =? 4+2=?

OpenStudy (anonymous):

9,6

OpenStudy (loser66):

dat sit